# Generative Semantic Communication for Joint Image Transmission and Segmentation

### v1

@@ V-A Experimental Setup
Baselines. For the image reconstruction task, we consider four baseline methods: 1) JPEG+LDPC+QAM(16/64), which adopts conventional separate source and channel coding and utilizes classical QAM for modulation; 2) wireless image transmission transformer (WITT) [20], which uses Swin-Transformer to extract semantic information; 3) Deep JSCC [21], in which the JSCC encoder consists of five convolutional layers while the JSCC decoder uses transposed convolution layers to upsample the compressed feature maps back to the images; 4) ADJSCC [22], which dynamically adjusts the coding strategy to optimize resource utilization, achieving efficient transmission.
For the image segmentation task, we adopt two baseline methods: 1) JPEG+LDPC+QAM(16/64)+FCN, which uses conventional separate source and channel coding, while utilizing a fully-connected network [23] to segment images at the receiver; 2) Deeplab JSCC [24], which adopts the Deeplab network as the fundamental architectures for the JSCC encoder and decoder. Through atrous convolution, it effectively captures the multi-scale contextual information to complete the image segmentation task.
@@
Model parameters. We conduct end-to-end training for all methods. In particular, we use AdamW as an optimizer, and the learning rate is set as 0.0001. The momentum coefficients are set as 0.5 and 0.999 with a weight decay of 0.01 to prevent overfitting. Other model parameters are listed in Table I.
@@ V-B Performance on Image Reconstruction
Fig. 4 and Fig. 5 show the performance comparison of the proposed framework with the other four baselines for both low-resolution and high-resolution images. In the low-resolution image case, our method consistently achieves the best reconstruction performance among all methods, particularly outperforming the ADJSCC and WITT. This result highlights the importance of semantic KBs in image reconstruction. In the high-resolution image case, our method performs slightly better than that of other baselines when SNR is low. This is because high-resolution images contain more texture information than low-resolution images. As a result, even minor noise or pixel loss will significantly affect the perceived quality of the reconstructed image. Therefore, the impact of noise is more pronounced on high-resolution images when SNR is low. However, the performance gaps between our method and other baselines increase with SNR, which demonstrates the effectiveness of the proposed method when the communication environment is fine. For better comparison, we visualize the reconstructed images of different methods when SNR = 18 dB, as shown in Fig. 6. It is observed that our method achieves the best visual quality with smooth facial details. More importantly, our framework achieves the lowest communication overhead among all methods.
Figure 4: PSNR v.s. SNR in low-resolution image case.
Figure 5: PSNR v.s. SNR in high-resolution image case.
@@ V-C Performance on Image Segmentation
Fig. 7 shows the image segmentation performance of all methods. For a fair comparison, the communication overhead of all methods is bounded by 20 Kbits. It is seen that under the same communication overhead, our method achieves the highest image segmentation accuracy compared to other baseline methods. In particular, when the SNR is greater than 0dB, the performance of our method will stabilize, demonstrating its robustness. Moreover, for the LDPC+JPEG+QAM methods, when the SNR is low ($<$12dB for 16QAM or $<$18dB for 64QAM), the receiver cannot accurately reconstruct the original image. Therefore, the FCN model cannot perform effective image segmentation, such that the image segmentation accuracy under low SNR is omitted. The bottom half of Fig. 7 provides the visualization results of all methods. It is seen that the segmentation result of our method is consistent with the ground-truth label, which not only identifies the target objects but also preserves the edge details. In contrast, the Deeplab JSCC cannot accurately distinguish the target object, while the boundaries are blurred. The traditional LDPC+JPEG+QAM methods perform even worse with erroneous areas in the segmentation results.
@@ VI Conclusion
In this paper, we propose a generative semantic communication system for both image transmission and segmentation. Two generative AI schemes of Swin-Transformer and diffusion model are adopted to build the semantic KBs and JSCC decoder, significantly reducing the data traffic while improving the communication efficiency. Numerical results show that our method can achieve better performance than several baseline methods. Our designs reveal the potential of using generative AI schemes for simultaneously handling multiple tasks and provide a new direction for achieving multi-task generative semantic communication.

### v2

@@ V-A Experimental Setup
Baselines. For the image reconstruction task, we consider four baseline methods: 1) JPEG+LDPC+QAM(16/64), which adopts conventional separate source and channel coding and utilizes classical QAM for modulation; 2) wireless image transmission transformer (WITT) [18], which uses Swin-Transformer to extract semantic information for transmission; 3) Deep JSCC [19], in which the JSCC encoder consists of five convolutional layers while the JSCC decoder uses transposed convolution layers to reconstruct the image; 4) ADJSCC [20], which dynamically adjusts the coding strategy to optimize resource utilization, achieving efficient transmission.
For the image segmentation task, we adopt two baseline methods: 1) JPEG+LDPC+QAM(16/64)+FCN, which uses conventional separate source and channel coding, while utilizing a fully-connected network [21] to segment images at the receiver; 2) Deeplab JSCC [22], which adopts the Deeplab network as the fundamental architecture for the JSCC encoder and decoder. Through atrous convolution, it captures the multi-scale contextual information to achieve effective image segmentation.
@@
Model Parameters. We perform end-to-end training for all methods. Specifically, we use AdamW as the optimizer with a learning rate of 0.0001 to control the step size during optimization. Momentum coefficients are set to 0.5 and 0.999 to accelerate convergence and reduce oscillations. A weight decay of 0.01 is applied to prevent overfitting. Other model parameters are listed in Table I.
@@ V-B Performance on Image Reconstruction
Fig. 4 shows the performance comparison of the proposed method with the four baselines for both low-resolution and high-resolution images. In the low-resolution image case, our method consistently achieves the best reconstruction performance among all methods, particularly outperforming the ADJSCC and WITT. This result highlights the importance of semantic KBs in image reconstruction. In the high-resolution image case, our method performs slightly better than other baselines when SNR is low. This is because high-resolution images contain more texture information than low-resolution images. As a result, even minor noise or pixel loss will significantly affect the perceived quality of the reconstructed image. Therefore, the impact of noise is more pronounced on high-resolution images when SNR is low. However, the performance gaps between our method and other baselines increase with SNR, which demonstrates the effectiveness of the proposed method when the communication environment is fine. For a better comparison, we visualize the reconstructed images of different methods when SNR = 18 dB, as shown in Fig. 5. It is observed that our method achieves the best visual quality with smooth facial details. In contrast, the visual results of WITT and ADJSCC exhibit noticeable color deviations. It is because both methods prioritize restoring structural information, such as edges and contours, at the expense of color information preservation. It is worth mentioning that our method also achieves the lowest communication overhead among all methods.
Figure 4: PSNR v.s. SNR in (a) low-resolution image case; (b) high-resolution image case.
@@ V-C Performance on Image Segmentation
Fig. 6 shows the image segmentation performance of all methods. For a fair comparison, the communication overhead of all methods is bounded by 20 Kbits. It is observed that under the same communication overhead, our method always achieves higher image segmentation accuracy compared to other baseline methods. In particular, when the SNR is greater than 0 dB, the performance of our method will stabilize, demonstrating its robustness. Moreover, for traditional communication methods (LDPC+JPEG+QAM(16/64)+FCN), when the SNR is low ($<$12dB for 16QAM or $<$18dB for 64QAM), the receiver cannot reconstruct the original image and the FCN model cannot perform effective image segmentation. Hence, the image segmentation accuracy under low SNR is omitted. The bottom half of Fig. 6 provides the visualization results of all methods. It is seen that the segmentation result of our method is consistent with the ground-truth label, which not only identifies the target objects but also preserves the edge details. In contrast, the Deeplab JSCC cannot accurately distinguish the target objects while the boundaries are blurred. Besides, the traditional communication methods perform even worse with erroneous areas in the segmentation results.
@@ VI Conclusion
In this paper, we propose a generative semantic communication system for both image reconstruction and segmentation tasks. Two generative AI schemes of Swin-Transformer and diffusion model are adopted to build the semantic KBs and JSCC decoder, significantly reducing the communication overhead while improving the transmission efficiency. Experimental results show that our method can achieve better performance than several baseline methods. Our designs reveal the potential of using generative AI schemes to handle multiple tasks and provide a new direction for achieving multi-task generative semantic communication.
Acknowledgement
This work was supported in part by NSFC with Grant No. 62293482, the Basic Research Project No. HZQB-KCZYZ-2021067 of Hetao Shenzhen-HK S&T Cooperation Zone, the Shenzhen Outstanding Talents Training Fund 202002, the Guangdong Research Projects No. 2017ZT07X152 and No. 2019CX01X104, the Guangdong Provincial Key Laboratory of Future Networks of Intelligence (Grant No. 2022B1212010001), and the Shenzhen Key Laboratory of Big Data and Artificial Intelligence (Grant No. ZDSYS201707251409055).